OpenStudy (anonymous):

the series sum (-1)^n / ln(n+1) is divergent? conditionally convergent? absolutely convergent?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ (-1)^n }{ \ln (n+1) }\]

OpenStudy (anonymous):

this should be alternating type

OpenStudy (anonymous):

isn't this absolutely convergent?

OpenStudy (anonymous):

my note says ratio,root,absolute value, alternating series are absolutely convergent, but i don't know why..

OpenStudy (anonymous):

the correct answer is conditionally convergent, if you take the absolute value of the series, you will get\[\sum_{n=1}^{\infty}\frac{1}{\ln(n+1)}\]try to test this series with the comparison test, compare with 1/n

OpenStudy (anonymous):

incorrect you will compare this with 1/ n+1 and not 1/n

OpenStudy (anonymous):

if you want to approximate , we can say 2.3 *(n+1) = ln (n+1), pull out 2.3 as constant and we are left with n+1

OpenStudy (anonymous):

but the overall answer is correct

OpenStudy (anonymous):

why is it incorrect? the inequality\[\ln (n + 1) < n\]is true for n > 0, which implies\[\frac{1}{\ln(n + 1)} > \frac{1}{n}\]for n > 0